Home‎ > ‎Crew 131 - Embry Riddle‎ > ‎

1212-CommandersReport

Commander Report
12/12/2013
Chelsea Iwig

Today was a much needed slow day for Crew 131. After all of the stress that the crew was under 
throughout the day yesterday, they needed the opportunity to recover and refocus themselves on 
making it through the remainder of their time on Mars. The morning began with one of the crew 
science officers rising to cook breakfast for the rest of the crew. As the smell of bacon and eggs 
wafted through the second floor of the Hab, the other crewmembers began emerging from their rooms 
with their stomachs growling from the smell of the delicious food being prepared. The crew enjoyed 
their breakfast burritos and then went back into their rooms to finish their daily preparations. 
    After receiving approval for the EVA planned for the day, Team 2 began suiting up to go 
explore a mountain 4km north of the Hab. Since the crew was finally beginning to recover from the 
first high intensity workout, two of the team members decided to grab their baseball gear and go 
out to play the first game of catch on the red planet before heading out on the Mars Rovers for 
their EVA. Much like when Al Shepherd became the first person to hit a golf ball on the lunar 
surface during the Apollo 14 mission, Crew 131 thought it would be a fun activity to bring baseball 
to the Red Planet. However, after finishing their little game of catch and preparing to head out, 
rover 3 shut down and they were unable to get it back up and running after several minutes of 
trying. Therefore the decision was made that two of the teammembers would continue with the planned 
EVA while the Commander remained at the Hab. This enabled two other crew members to leave in the 
Explorer capsule to head back to the nearby service station to pick up a package and the equipment 
needed to run the Mars Observatory. Therefore, with two crewmembers out on EVA and two in the 
Explorer headed for the service station, the remaining two crewmembers stayed behind to clean 
dishes and cook a late lunch. However, to their surprise, the two members of EVA team 2 returned a 
short time later as they had decided that they did not want to continue on with their EVA without 
their Commander. Instead, they went back to the scene of yesterday’s accident to pick up any pieces 
of helmet that had not been picked up by Team 1 before they returned to the Hab. 
    Once Crew 131 was all united back in the Hab again, they were able to relax for the rest of 
the afternoon and just enjoy each other’s company before beginning their work on the evening’s 
Communication Reports. Tonight, after submitting their reports and completing the surveys for their 
research, the crew plans on making some popcorn and watching Star Wars the Empire Strikes Back 
together to close out a nice and relaxing day on the Martian surface. 

----------------------------

CommanderCheckIn Report
12/12/2013
Chelsea Iwig
Crew Physical Health:
All crewmembers are in good health
   
Eva Depart Time:
1141
Eva Return Time:
1315
Narrative Of Field Mission Results:
EVA team two was supposed to have their first steep walking EVA today, but to get to their destination, they needed to take the ATVs approximately 4km north of the Hab. The team suited up and all three ATVs were running and ready to go. However, once the crew finished packing up the ATVs, one of the ATVs died. The crew tried to get it running again, but they were unsuccessful. The Commander made the decision to let the other two crewmembers carry on with the EVA as normal while she stayed behind to get some work done in the Hab. Even so the other two team members decided, rather than to go all the way to their planned destination, to just go back to the scene of the ATV accident yesterday and pick up the pieces of EVA helmets that were left behind. They then returned to the Hab early as they felt it would not be right to continue the EVA without their third teammate. 
Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Eva Depart Time:

Eva Return Time:

Narrative Of Field Mission Results:

Hab Maint:
   The pipes were frozen again this morning, so the Crew Engineer thawed out the pipes this morning. She also continues to turn the water pump on every two hours during the day and the night just to make sure it doesn’t freeze. 
ATV 4’s battery was re-installed after being charged, but the ATV still doesn’t work. 
ATV 3 needs to be worked on as it shut down this morning after running for several minutes and we could not get it started again. 
The EVA packs were worked on last night to make sure we would have at least three working suits that match up with helmets. Currently, suits 1, 3, and 4 are the only ones that are working optimally. The others work somewhat, but the airflow is very weak. 
Rep Sched:
   Commander Check-In, Commander Report, Engineering Report, HSO Report, EVA Report, EVA Plan, Journalist Report, Science Report
Tomorrow Plan:
  Tomorrow, the crew will begin the day bright and early to complete the second high intensity workout for the exercise study. Once the crew has had a few minutes to recover, EVA team 1 will go out on a low intensity walking EVA to test out the new EVA suits. This testing has been saved for our first low intensity walking EVA as the new helmets are much more fragile than the old helmets and we don’t want to risk hitting them on a rock while climbing or them falling off of the ATVs and breaking. Also, during this EVA, the GPS armbands will again be tested. Once the EVA team returns, two crewmembers will go out to the observatory to work on getting the telescope set up now that the crew has been able to retrieve the necessary telescope pieces during their trip to pick up a package today from the Hollow Mountain Service Station. Meanwhile, two other crewmembers will be conducting another session of the space suit glove study and the plants for the aeroponics study will be maintained. In the evening, the crew will complete reports, make dinner, clean up the Hab, and then go out to the observatory to complete some star observations.  
Inventory:
   Inventory of the HSO cabinet was conducted late last night and posted on the door of the cabinet. 
Support Req:
   New filters for the EVA packs are needed. 
Miscelleneous:
   There are still at least two mice and now we have seen 3 very large rats on the loose in the Hab. More traps have been set and our Crew Engineer has moved down from the loft into the remaining open stateroom as she saw the other two rats sitting very close to her bed while she was sleeping last night.